The graphs of functions f, g are given. Find the values ​​of (a) f (−4) and f (3).
What value of (b) has f (x) = g (x)?
(c) Find the approximate solution of the equation f (x) = −1.
(d) In what interval does f decrease?
(e) Find the determinable range of the function f and the range of values.
(f) Find the range and the range of values ​​of the function g.
1
Expert's answer
2020-03-16T12:02:05-0400

The graphs of functions f, g are given.



(a)Find the values ​​of "f(-4)" and "f(3)."


"f(-4)=1, f(3)=0.8"

(b) What value of "x" has "f(x)=g(x)"


"x_1=-6, f(-6)=5=g(-6)"

"x_2=3, f(3)=1=g(3)"

"x_3=4, f(4)=0=g(4)"

(c) Find the approximate solution of the equation "f(x)=-1"


"x\approx4.7"

(d) In what interval does "f" decrease? 

"f" decreases for "x\in (-\infin,-2)\cup(2,\infin)"


(e) Find the determinable range of the function f and the range of values. 

the determinable range of the function "f: (-\infin,\infin)"

the range of values":[-4,5]"


(f) Find the range and the range of values ​​of the function g.

the range of the function "g: (-\infin,\infin)"

the range of values ":[-1,5]"